Nick's telephone company charges $0.06 per minute for weekend calls and
$0.08 per minute for calls made on weekdays. This month Nick was billed for
544 minutes. The charge for these minutes was $41.84.  How many minutes did he talk on weekends and how many minutes did he talk on​ weekdays?

Answer by Boreal(15235)

use 6, 8 and 4184
W+D=544
6W+8D=4184
-6W-6D=-3264
2D=920
D=460@8 cents or $36.80
W=84@6 cents or $5.04
Total $41.84